a) Define viscosity.



b) A matter has a body-centered cubic unit cell with a length of 330 π‘π‘š along an edge and a density of 8.57 𝑔 π‘π‘šβˆ’3 . Calculate the atomic mass of the matter.



c) Find the total energy of a moving object of mass 6.3 π‘˜π‘” and momentum 11.5 Γ— 106 π‘˜π‘”π‘šπ‘  βˆ’1 .




1
Expert's answer
2021-12-09T09:22:44-0500

a) TheΒ viscosityΒ of a fluid is a measure of its resistance to deformation at a given rate.


b)


"(3.3\cdot10^{-10})^2=\frac{2M}{8.57\cdot6.023\cdot10^{23}}\\M=93"

c)


"E=\frac{p^2}{2m}\\E=\frac{(11.5\cdot10^6)^2}{2(6.3)}=1.05\cdot10^{13}\ J"